# Hybrid Quantum-Classical Recurrent Neural Networks

### v1

@@ Appendix C Experimental Settings and Test Accuracy Statistics Across Runs
Table 6: Accuracy statistics on IMDB test set across 100 runs for each nonlinearity variant. Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp. Among all tasks, IMDB showed the greatest variability in QRNN performance across random seeds in development. This behavior may align with known sensitivities in training variational PQCs [Grant et al., 2019]. We therefore also report stats where we remove failed runs ($70\%$ accuracy, well below simple baselines such as BoW), indicated by *. For the three nonlinearities 40, 42 and 25 failed runs were observed each. The results also indicate that GELU nonlinearity reduces the sensitivity compared with the other two.
@@
Table 7: Accuracy statistics on MNIST and pMNIST test sets across 50 runs for each nonlinearity variant. Qubit count $q$ and total measurements $m$; embedding dimension $e$; parameter count pp.
@@
Table 8: BLEU evaluations on the Multi30K German to English test set across 20 runs for each nonlinearity variant. Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp.
@@ Appendix D Attention Alignments
We observe that the hybrid model can manage spatial and syntactic shifts while capturing clause-level structure and semantics through its measurement-driven hidden states and soft attention as well as the LSTM baseline. It is evident that the model handles compound verb constructions and semantic expansion, in sentences like “Diese Band bereitet sich auf einen Auftritt vor Publikum in einer Kirche vor” (Fig. 4(a)) and “Zwei grün gekleidete Männer bereiten in einem Restaurant Essen zu” (Fig. 4(b)), where German separable verbs—“bereitet … vor” and “bereiten … zu”—are correctly reconstructed into the English verb phrases “is preparing to perform” and “preparing”, respectively. The soft attention allowed the model to attend across non-contiguous source tokens, enabling reassembly of verb phrases. Additionally, lexical expansions such as “Publikum” $\rightarrow$ “a crowd of people” (Fig. 4(a)) and “gekleidete Männer” $\rightarrow$ “men in green outfits” (Fig. 4(b)) demonstrate contextually appropriate semantic elaboration beyond literal translation.

### v2

@@ Appendix C Experimental Settings and Test Accuracy Statistics Across Runs
Table 6: Accuracy statistics on IMDB test set across 100 runs for each nonlinearity variant. Qubit count $q$, total readouts $m$; embedding dimension $e$; parameter count pp. Among all tasks, IMDB showed the greatest variability in QRNN performance across random seeds in development. This behavior may align with known sensitivities in training variational PQCs [Grant et al., 2019]. We therefore also report stats where we remove failed runs ($70\%$ accuracy, well below simple baselines such as BoW), indicated by *. For the three nonlinearities 40, 42 and 25 failed runs were observed each. The results also indicate that GELU nonlinearity reduces the sensitivity compared with the other two.
@@
Table 7: Accuracy statistics on MNIST and pMNIST test sets across 50 runs for each nonlinearity variant. Qubit count $q$ and total readouts $m$; embedding dimension $e$; parameter count pp.
@@
Table 8: BLEU evaluations on the Multi30K German to English test set across 20 runs for each nonlinearity variant. Qubit count $q$, total readouts $m$; embedding dimension $e$; parameter count pp.
@@ Appendix D Attention Alignments
We observe that the hybrid model can manage spatial and syntactic shifts while capturing clause-level structure and semantics through its readout-based hidden states and soft attention as well as the LSTM baseline. It is evident that the model handles compound verb constructions and semantic expansion, in sentences like “Diese Band bereitet sich auf einen Auftritt vor Publikum in einer Kirche vor” (Fig. 4(a)) and “Zwei grün gekleidete Männer bereiten in einem Restaurant Essen zu” (Fig. 4(b)), where German separable verbs—“bereitet … vor” and “bereiten … zu”—are correctly reconstructed into the English verb phrases “is preparing to perform” and “preparing”, respectively. The soft attention allowed the model to attend across non-contiguous source tokens, enabling reassembly of verb phrases. Additionally, lexical expansions such as “Publikum” $\rightarrow$ “a crowd of people” (Fig. 4(a)) and “gekleidete Männer” $\rightarrow$ “men in green outfits” (Fig. 4(b)) demonstrate contextually appropriate semantic elaboration beyond literal translation.